OpenStudy (anonymous):

Please help!! fan and medal How can 1/4x− 3 = 1/2x+ 8 be set up as a system of equations? 4y + 4x = −12 2y + 2x = 16 4y − x = −12 2y − x = 16 4y + x = −12 2y + x = 16 4y − 4x = −12 2y − 2x = 16

OpenStudy (anonymous):

@Hero

hero (hero):

We can start with y = (1/4)x - 3 y = (1/2)x + 8 Now multiply both sides of the first equation by 4, then multiply both sides of the second equation by 2. What do you get afterwards?

OpenStudy (anonymous):

well I did the whole process and got 14. but I don't understand how to get the format of the choices above

OpenStudy (anonymous):

Would it be 'B' ?

OpenStudy (anonymous):

wow this is a really idiotic question if you don't mind me saying so who on earth would want to solve this simple linear equality by turning it in to a system of equations??

OpenStudy (anonymous):

\[y=\frac{1}{4}x-3\\4y=x-12\\ 4y-x=-12\]

OpenStudy (anonymous):

forget the second equation, that leaves you only one choice, what you said B

OpenStudy (anonymous):

Thank you (: